An object with a density of 28 g/cm squared has a mass of 17.5 grams. What is the volume?
1
Expert's answer
2018-08-30T03:52:03-0400
Solution:

m= p*V, V =m/p=17.5/28 = 0.625 cm^3

Answer: V = 0.625 cm^3